OpenStudy (anonymous):

find the determinant of each matrix. show work -2,4,1 3,0,-1 1,2,1

OpenStudy (anonymous):

I'll use a cofactor expansion along the second row: \[\begin{align*}\begin{vmatrix}-2&4&1\\3&0&1\\1&2&1\end{vmatrix}&=(-3)\begin{vmatrix}4&1\\2&1\end{vmatrix}+0\begin{vmatrix}-2&1\\1&1\end{vmatrix}+(-1)\begin{vmatrix}-2&4\\1&2\end{vmatrix}\\ &=-3\begin{vmatrix}4&1\\2&1\end{vmatrix}-\begin{vmatrix}-2&4\\1&2\end{vmatrix}\\ &=-3(4-2)-(-4-4)\\ &=-6-(-8)\\ &=2 \end{align*}\]

OpenStudy (anonymous):

wait how come you put a -3 up top when its a postitive?

OpenStudy (anonymous):

Here's the pattern when doing a cofactor expansion: \[\begin{bmatrix}+&-&+&\cdots\\ -&+&-&\cdots\\ +&-&+&\cdots\\ \vdots&\vdots&\vdots&\ddots\end{bmatrix}\] Basically, what this means is that you either put a plus or minus sign on a given number, depending on where it is in the given matrix, and multiply that by the minor matrix that's left.